A certain flight arrives on time 80 percent of the time. Suppose 174 flights are randomly selected. Use the normal approximation to the binomial to approximate the probability that

​(a) exactly 148 flights are on time.

​(b) at least 148 flights are on time.

​(c) fewer than 133 flights are on time.

​(d) between 133 and 150​, inclusive are on time.
Answer #1

Let X denote the number of flights arriving on time. Then X\sim Bin(n=174,p=0.80)

Using Normal approximation, X\sim N(\mu=np=139.2,\sigma^2=np(1-p)=27.84)

a)

Required probability = P(X=148)=P(147.5\leq X\leq 148.5)

{Due to continuity correction as a result of using Normal approximation for Binomial distribution}

=P(\frac{147.5-139.2}{\sqrt{27.84}}<Z<\frac{148.5-139.2}{\sqrt{27.84}})=P(1.57<Z<1.76)

=P(Z<1.76)-P(Z<1.57)=0.96080-0.94179=0.01901

b)

Required probability = P(X\geq 148)

=P(Z>\frac{148-139.2}{\sqrt{27.84}})=P(Z>1.67)=1-P(Z<1.67)

=1-0.95254=0.04746

c)

Required probability = P(X<133)=P(X\leq 132.5)

{Due to continuity correction as a result of using Normal approximation for Binomial distribution}

=P(Z<\frac{132.5-139.2}{\sqrt{27.84}})=P(Z<-1.27)=P(Z>1.27)

=1-P(Z<1.27)=1-0.89796=0.10204

d)

Required probability = P(133\leq X\leq 150)

=P(\frac{133-139.2}{\sqrt{27.84}}<Z<\frac{150-139.2}{\sqrt{27.84}})=P(-1.18<Z<2.05)

=P(Z<2.05)-P(Z<-1.18)=P(Z<2.05)-P(Z>1.18)

=P(Z<2.05)-[1-P(Z<1.18)]

=0.97982-[1-0.88100]=0.86082